[Hallc_running] [New Logentry] RC Daily Update

jones at jlab.org jones at jlab.org
Sat Sep 29 10:00:08 EDT 2018

Logentry Text:
<p>Below is info from yesterday's RC meeting. We finished up the optics check out of the HMS/SHMS at around 4:30 yesterday.</p><p>Started the LH2 elastic data taken. Since the pass change, beam has been unstable and we have not been able to</p><p>get reliable stable high current running. As of 9:00 on Sat, we have finished 4 of the  8 elastic single arm settings.</p><p> </p><p><strong>Main issue is with HMS Dipole LN2 level.</strong> https://logbooks.jlab.org/entry/3600022</p><p>The users have to manual change the valve for JT3. There is a alarm to let you know if it get outside its limits</p><p>but it is best to monitor it so that try one can keep it from reaching extremes.</p><p> </p><p> </p><p>Since we are still in the commissioning phase, we will meet this afternoon at 4:30pm</p><p>in the second floor counting house: <a href="https://hallcweb.jlab.org/wiki/index.php/RC_daily_meetings_and_daily_run">https://hallcweb.jlab.org/wiki/index.php/RC_daily_meetings_and
 _daily_run</a></p><p> </p><p>*''Status''</p><ol><li>**Yesterday the HMS quad work was completed by about 1pm and we started the optics study of HMS/SHMS at high momentum.</li><li>**Beam was bad until around 6pm when MCC had solved the problem of the many trips coming from Hall B beam.</li><li>** When changing to set the HMS at 6 GeV , The HMS Q3 would not stay at a given set current. This put into doubt the the higher momentum HMS optics data (taken at 6.6 and 6.3).</li><li>** MCC was able to deliver up to 60uA on the owl shift. Given problem with HMS, decided to do the BCM which was completed. https://logbooks.jlab.org/entry/3599568</li><li>**After BCM data was taken , then continued with the SHMS optics which was completed.</li><li>**Turned on EDTM pulser for BCM calib data. Left it on.</li><li>**The Hall went into controlled access around 8:30 for work on the HMS Q3. Found that the opto-isolator was not pushed in all the way.</li><li>**Around 9:00, MCC started the pass chang
 e for Hall A.</li><li>**Around noon, started!
  to receive beam and restarted the optics study for HMS ( repeated some SHMS). finsihed around 4:20.</li><li>**Updated the EPICS BCM names and calibrations https://logbooks.jlab.org/entry/3599765</li><li></li></ol><p><br>*''Reports''</p><ol><li>**SHMS Cerenkov https://logbooks.jlab.org/entry/3599771</li><li>** SHMS PID check https://logbooks.jlab.org/entry/3599464</li></ol><p><br>*''Issues''</p><ol><li>** Steve Wood to comment on importance of watching closely the GUI and TV screen when moving spectrometers.</li><li>** Changing the attenuator of the SHMS Cer sum that is used to monitor that leg of the trigger for 0.3 to 1.0 . https://logbooks.jlab.org/entry/3599768</li><li>** HMS UPS is bad. Jumpered interlock.  https://logbooks.jlab.org/entry/3599763</li><li>** Gas shed https://logbooks.jlab.org/entry/3599543</li></ol><p><br><br></p>


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3600543
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20180929/0118803d/attachment.html>

More information about the Hallc_running mailing list